You draw two slips of paper from a hat containing five slips, marked 1 through 5, with replacement. What is the probability that exactly one of the numbers is a five?
The following is my thought process: So for the first slip of paper, the chance of getting a 5 would be 1/5. Since the slip of paper can be replaced, the chance of getting another 5 would still be 1/5. But since you only want one 5, would the answer be 1/10?

Answer by ewatrrr(24785)

P(A or B) = P(A) + P(B) - P(A and B) = 1/5 + 1/5 - (1/5)(1/5) = 2/5 - 1/25